# Generative Modeling with Explicit Memory

### v1

@@ Abstract
Corresponding author.* Equal contribution. Peng is the project leader.
Recent studies indicate that the denoising process in deep generative diffusion models implicitly learns and memorizes semantic information from the data distribution. These findings suggest that capturing more complex data distributions requires larger neural networks, leading to a substantial increase in computational demands, which in turn become the primary bottleneck in both training and inference of diffusion models. To this end, we introduce Generative Modeling with Explicit Memory (GMem), leveraging an external memory bank in both training and sampling phases of diffusion models. This approach preserves semantic information from data distributions, reducing reliance on neural network capacity for learning and generalizing across diverse datasets. The results are significant: our GMem enhances both training, sampling efficiency, and generation quality. For instance, on ImageNet at $256\times 256$ resolution, GMem accelerates SiT training by over 46.7×46.7$\times$46.7 ×, achieving the performance of a SiT model trained for $7M$ steps in fewer than $150K$ steps. Compared to the most efficient existing method, REPA [70], GMem still offers a 16×16$\times$16 × speedup, attaining an FID score of 5.75 within $250K$ steps, whereas REPA requires over $4M$ steps. Additionally, our method achieves state-of-the-art generation quality, with an FID score of 3.56 without classifier-free guidance on ImageNet $256\times 256$. Our code is available at https://github.com/LINs-lab/GMem.
Figure 1: GMem significantly enhances the training and sampling efficiency of diffusion models on ImageNet $256\times 256$. We propose an external memory bank that enhances the semantic information storage, improving both training stability and sampling efficiency in diffusion models. Sub-figure (a) illustrates the key concept of GMem, where $\boldsymbol{\epsilon}$ represents the initial noise and $\mathbf{x}_{0}$ the generated samples. In GMem, we disentangle the generalization and memorization capabilities originally intertwined within diffusion models, assigning memorization to an external memory bank $\mathbf{M}$. This decoupling reduces computational and capacity overhead for memorizing internal semantic information, thereby accelerating the overall process. Sub-figure (b) demonstrates the training efficiency of GMem on ImageNet $256\times 256$. At an FID of $5.8$, GMem achieves over 16×16$\times$16 × speedup relative to REPA [70]. At an FID of $7.3$, GMem achieves over 46.7×46.7$\times$46.7 × speedup relative to SiT [44]. Sub-figure (c) illustrates the sampling efficiency of GMem on ImageNet $256\times 256$. At the same target level of FID, GMem requires 5×5$\times$5 × less NFEs compared to REPA and 10×10$\times$10 × less NFEs compared to SiT.
@@ 1 Introduction
Deep generative models based on denoising [68, 24, 59, 47, 9, 40, 8, 37]—which prioritize the generation of high-quality, realistic data—have achieved notable success within the deep learning community. These methods demonstrate exceptional performance in complex tasks such as zero-shot text-to-image and video generation [50, 56, 15, 51, 4]. However, training and sampling diffusion models suffer from the issue of high computational burden [28, 29], due to the necessity of using neural networks with larger capacities for better empirical performance [28, 29].
In the meanwhile, a line of research examines the diffusion models from the perspective of representation learning [36, 66, 8, 46], where diffusion models can capture semantic features within their hidden layers, with advanced models producing even stronger representations [66]. Yang and Wang [68] illustrate that diffusion models strike a balance between learning meaningful features and regularizing the model capacity. Kadkhodaie et al. [27] further interpret the generalization behavior in diffusion models, indicating that diffusion models aim to learn and encapsulate semantic information from the training data, subsequently generalizing to the true data distribution.
In this paper, we extend these insights [68, 27] and conjecture that the functions of diffusion models can be implicitly divided into two components: (i) memorization of semantic information and (ii) generalization to true data. The intrinsic generation procedure of diffusion model then can be explained as transforming input noisy samples into the internal semantic distribution of a model, followed by “decoding” it into the data distribution. Specifically,
@@
    in the training stage, diffusion models require significant neural network parameters and computational resources to extract and memorize semantic information from noisy samples;
@@
    in the sampling/inference stage, models often remap Gaussian noise to the internal semantic distribution before retrieving the data distribution, rather than directly transforming noise to data.
Building upon the aforementioned analyses, the efficiency challenges of training and sampling diffusion models stem from the need to learn and memorize the internal semantic information, resulting in higher computational cost and greater demands on model capacity. As a remedy, we propose GMem, a new paradigm that uses an external memory bank to accompany the diffusion models to approximate the true data distribution, therefore avoiding memorizing the data distributions in the neural network. Extensive experiments demonstrate the effectiveness and efficiency of our approach, on both diffusion transformers [44] and convolutional neural networks [24, 61, 28]. For example, integrating our GMem with SiT on ImageNet $256\times 256$ only requires $150K$ training iterations to achieve an FID of $7.3$, which is 46.7×46.7$\times$46.7 × faster compared to the baseline SiTs [44] and 16×16$\times$16 × than REPA [70]. Under the same number of training iterations as $400K$, our model enhances the sampling efficiency by 10×10$\times$10 ×, where our model requires only $25$ sampling steps to achieve an $\texttt{FID}=12.3$ while the vanilla SiT needs $250$ sampling steps. Ultimately, our approach achieves SoTA performance with an FID of $3.56$ after $2M$ training steps, without classifier-free guidance.
We outline the main contributions of this paper as follows:
@@
    We propose that diffusion models operate through two fundamental perspectives: memorization and generalization. Utilizing neural networks for memorization incurs significant computational and model capacity burdens.
@@
    We present GMem, a simple yet extremely effective approach to construct an external memory bank that stores semantic information inherent in data distributions, alleviating the memorization burden on neural networks within the diffusion models.
@@
    Our framework significantly enhances efficiency and quality of the SoTA diffusion models, for both training and inference (see Figure 1 and Figure 2).
Figure 2: Selected samples on ImageNet $256\times 256$. This figure showcases a selection of images generated by GMem. We utilized the SiT-XL/2 backbone and trained GMem for $2M$ steps without classifier-free guidance.
@@ 2 Related work
Generative models—including Generative Adversarial Networks (GANs) [19, 57, 67], Variational Autoencoders (VAEs) [32, 21], flow-based methods, and diffusion-based methods [24, 12, 45]—aim to learn the data distribution $p(\mathbf{x})$ and generate data through sampling, demonstrating remarkable performance in producing realistic samples [38]. Recently, diffusion-based methods utilize stochastic interpolation to model the forward process and subsequently convert the Gaussian distribution back into the original image distribution during the reverse process, yielding realistic samples. Such methods have achieved SoTA results in deep generative modeling and are the focus of this study [45, 60, 14].
Training diffusion-based models is computationally intensive and unstable [70, 60]. Additionally, the generative process necessitates multi-step sampling, incurring substantial computational costs [42] Accelerating the training and sampling processes of diffusion models has emerged as a significant challenge, prompting various efforts in previous studies to address this issue.
@@
To overcome the instability and computational inefficiency of diffusion models, recent studies  [70, 17, 45] have turned to representation learning to enhance diffusion models. On one hand, diffusion models are capable of learning high-quality representations [70]. For instance, Tang et al. [63] demonstrated that feature maps extracted from diffusion networks can establish correspondences between real images, indicating a strong alignment between the learned representations with actual image. Furthermore, Yang and Wang [68] conducted a detailed analysis of the trade-off between the quality of learned representations and the penalization of the optimal parameter spectrum.
@@
Conversely, well-trained representation models can improve performance and expedite the training of diffusion models. [45] accomplish this by adjusting the weighting function in the denoising score matching objective to enhance representation learning. REPA [70] introduces an alignment loss for intermediate layer representations, significantly accelerating the training process by over $17.5$ times. In general, these methods enhance diffusion modeling by employing regularization techniques, which implicitly integrate semantic knowledge from pre-trained representation models.
Self-supervised learning in diffusion models.
Self-supervised learning, as a powerful approach within representation learning that leverages unlabeled data to learn useful representations, has garnered increasing attention in the diffusion modeling community. Many recent works [39, 26, 2] have adopted self-supervised learning techniques to enhance the representation learning capabilities of diffusion models, thereby improving their performance and efficiency. In particular, Li et al. [39] proposed a novel diffusion training paradigm that introduces an additional diffusion model to generate a 1D feature map as a condition to guide the training of the generative model. However, the inclusion of an additional diffusion model introduces extra sampling during inference, which increases computational overhead and can introduce additional errors. In this work, differing from Li et al. [39], we directly utilize representation from ‘memory bank’ as condition. This straightforward strategy eliminates the need for an additional diffusion model, thereby enhancing representation learning while reducing computational costs. It significantly improves both the training speed and the effectiveness of the model. In this work, differing from Li et al. [39], we directly utilize representation from ‘memory bank’ as condition. This straightforward strategy not only eliminates the unnecessary computational burden but also serves as an effective self-supervised learning method that enabling the model to learn better representations. As a result, our method significantly enhances both the training speed and effectiveness of diffusion model.
Text to image generation.
Text-to-image generation seeks to synthesize realistic images from free-form textual prompts. Despite advancements, current text-to-image models continue to experience challenges with text-image misalignment [53, 35]. Recent works [6, 16] have attempted to address this issue by intervening in cross-modal attention activations that influence the internal states of Stable Diffusion models. However, this approach increases computational complexity and may impede the learning of optimal representations. Indeed, these models often require several days to weeks for training.
@@ 3 Preliminary
In this section, we provide a concise introduction to the training and sampling processes of flow-based and diffusion-based models. See Appendix G for more details.
Both diffusion-based and flow-based models [44] derive their training procedures from a deterministic $T$-step noising process applied to the original data, formalized as:
@@
According to (1), each marginal probability density $p_{t}(\mathbf{x}_{t})$ corresponds to the distribution of a Probability Flow Ordinary Differential Equation [61] (PF ODE) with velocity field $\mathbf{v}(\mathbf{x},t)$, defined as:
@@
Alternatively, the aforementioned noise-adding process can be formalized as a Stochastic Differential Equation [61, 59] (SDE):
@@
where $\mathbf{W}_{t}$ is a Wiener process [23], $\boldsymbol{m}(\mathbf{x}_{t},t)$ is the drift coefficient defined as $\boldsymbol{m}(\mathbf{x}_{t},t)=-\frac{1}{2}\beta(t)\mathbf{x}_{t}$, and $g(t)$ is the diffusion coefficient, set as $g(t)=\sqrt{\beta(t)}$ with $\beta(t)$ being a time-dependent positive function controlling the noise schedule.
@@
By solving the reverse-time SDE (4), starting from the initial state $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we can obtain $p_{0}(\mathbf{x}_{0})$, thereby estimating the true data distribution $p(\mathbf{x})$.
@@ 4 Methodology
procedure Train GMem($\mathbf{v}_{\boldsymbol{\theta}},\mathcal{D},\mathbf{M},T,\alpha_{t},\sigma_{t}$)
     Initialize model parameters $\boldsymbol{\theta}$
     for each training iteration do
         Sample a batch of data $\mathbf{x}_{0}\sim\mathcal{D}$
         Sample timesteps $t\sim\{0,\dots,T\}$ uniformly
         Generate noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
         Compute noisy data $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
         Sample memory snippet $\mathbf{s}\sim\mathbf{M}$ using (9)
         Predict velocity $\mathbf{v}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{s})$
         Compute loss using (8) and (10)
         Backpropagate and update $\boldsymbol{\theta}$ using Optimizer
     end for
end procedure
The motivation for our method arises from the observation that diffusion models inherently encode two distinct capabilities: generalization and memorization of semantic information within data [70, 68, 66, 8]. We propose a paradigm in diffusion models that separates generalization and memorization: neural networks are employed for generalization, while an external memory bank is used to store semantic information of data.
Specifically, we find that conventional generative models [24, 60] typically use a neural network to achieve both generalization and memorization of true data distributions. However, the UNet and Transformer-based neural networks [24, 49] used in modern diffusion models have limited parameters. This makes it challenging for these networks to memorize distribution information, leading to two main issues [27]: (a) limited parameters reduce the ability to memorize extensive data distributions, and (b) memorization via parameter optimization incurs high computational and time costs.
To address these challenges and alleviate the memorization burden on neural networks in diffusion models, we propose employing an external memory bank to serve as the memorization function, therefore significantly improving training efficiency and the models’ ability to capture distributions. The framework is depicted in Figure 3.

### v2

# GMem: A Modular Approach for Ultra-Efficient Generative Models

@@ Abstract
Recent studies indicate that the denoising process in deep generative diffusion models implicitly learns and memorizes semantic information from the data distribution. These findings suggest that capturing more complex data distributions requires larger neural networks, leading to a substantial increase in computational demands, which in turn become the primary bottleneck in both training and inference of diffusion models. To this end, we introduce GMem: A Modular Approach for Ultra-Efficient Generative Models. Our approach GMem decouples the memory capacity from model and implements it as a separate, immutable memory set that preserves the essential semantic information in the data. The results are significant: GMem enhances both training, sampling efficiency, and diversity generation. This design on one hand reduces the reliance on network for memorize complex data distribution and thus enhancing both training and sampling efficiency. On ImageNet at $256\times 256$ resolution, GMem achieves a 50×50$\times$50 × training speedup compared to SiT, reaching FID $=7.66$ in fewer than $28$ epochs ($\sim$4444 hours training time) [^1] All training time measurements are obtained on an 8×8$\times$8 ×H800 GPU cluster., while SiT requires $1400$ epochs. Without classifier-free guidance, GMem achieves state-of-the-art (SoTA) performance FID $=1.53$ in $160$ epochs with only $\sim 20$ hours of training, outperforming LightningDiT which requires $800$ epochs and $\sim$95959595 hours to attain FID $=2.17$.
Additionally, the memory bank supports training-free adaptation to new images not present in the training set. Our code is available at https://github.com/LINs-lab/GMem.
Machine Learning, ICML
[…]
Figure 1: GMem Significantly enhances training and sampling efficiency of diffusion models on ImageNet $256\times 256$. We propose decoupling memorization capabilities from the model by implementing a separate, immutable memory bank that preserves essential data information. Sub-figure (a) highlights the core concept of GMem, where $\boldsymbol{\epsilon}$ denotes input noise and $\mathbf{x}_{0}$ represents generated samples. In GMem, we disentangle generalization and memorization capabilities, assigning memorization to an external memory bank $\mathbf{M}$. This decoupling reduces computational and capacity overhead, thus accelerating the process. Sub-figure (b) demonstrates the training efficiency of GMem on ImageNet $256\times 256$. At an FID $=4.86$, GMem achieves over 25×25$\times$25 × speedup compared to REPA (Yu et al., 2024). At an FID $=7.66$, it achieves over 50×50$\times$50 × speedup relative to SiT (Ma et al., 2024). Sub-figure (c) illustrates sampling efficiency. At the same FID target, GMem requires 5×5$\times$5 × fewer NFEs compared to REPA and 10×10$\times$10 × fewer NFEs compared to SiT.
@@ 1 Introduction
Deep generative models based on denoising (Yang & Wang, 2023; Ho et al., 2020; Song et al., 2020a; Nichol & Dhariwal, 2021; Choi et al., 2021; Li et al., 2024; Chen et al., 2024c; Li et al., 2023c)—which prioritize the generation of high-quality, realistic data—have achieved notable success within the deep learning community. These methods demonstrate exceptional performance in complex tasks such as zero-shot text-to-image and video generation (Podell et al., 2023; Saharia et al., 2022; Esser et al., 2024; Polyak et al., 2024; Brooks et al., 2024). However, training and sampling diffusion models suffer from the issue of high computational burden (Karras et al., 2022, 2024), due to the necessity of using neural networks with larger capacities for better empirical performance (Karras et al., 2022, 2024).
In the meanwhile, a line of research examines the diffusion models from the perspective of representation learning (Li et al., 2023a; Xiang et al., 2023; Chen et al., 2024c; Mukhopadhyay et al., 2021), where diffusion models can capture semantic features within their hidden layers, with advanced models producing even stronger representations (Xiang et al., 2023). Yang & Wang (2023) illustrate that diffusion models strike a balance between learning meaningful features and regularizing the model capacity. Kadkhodaie et al. (2023) further interpret the generalization behavior in diffusion models, indicating that diffusion models aim to learn and encapsulate semantic information from the training data, subsequently generalizing to the true data distribution.
In this paper, we build upon the insights from prior works (Yang & Wang, 2023; Kadkhodaie et al., 2023) and propose a novel conjecture: the functionality of diffusion models can be implicitly decomposed into two distinct components: (i) memorization of semantic information and (ii) generalization to the true data distribution. We further posit that the intrinsic generation process of diffusion models can be interpreted as a transformation of noisy input samples into the model’s internal semantic representation, which is subsequently “decoded” into the target data distribution. To elaborate,
@@
    During the training stage, diffusion models rely on substantial neural network capacity and computational resources to extract and memorize semantic information from noisy samples;
@@
    During the sampling/inference stage, the models typically remap Gaussian noise to the internal semantic distribution before reconstructing the data distribution, rather than directly transforming noise into data.
Building upon the aforementioned analyses, the efficiency challenges of training and sampling diffusion models stem from the need to learn and memorize internal semantic information. This process inherently incurs high computational costs and imposes significant demands on model capacity.
As a remedy, we propose GMem, a new paradigm that decouples the memorization capability from the neural network by storing semantic information in an external memory bank. This separation eliminates the need for the neural network itself to memorize complex data distributions, enabling significant improvements: (1) accelerated training and sampling by avoiding the direct fitting of intricate semantic distributions; (2) flexible adjustment of memorization capacity in a training-free manner; (3) enhanced generalization, allowing the generation of images beyond the training set; and (4) reduced storage overhead by offloading memorization to the memory bank. Together, GMem provides a scalable and efficient framework for generative tasks.
Extensive experiments across diverse diffusion backbones (Ma et al., 2024; Yao & Wang, 2025) and image tokenizers (Yao & Wang, 2025; Chen et al., 2024a) validate the effectiveness and efficiency of GMem. On ImageNet $256\times 256$, GMem achieves SoTA performance FID $=1.53$ in $160$ epochs with only $\sim$20202020 hours training time without classifier-free guidance.
Under the same number of training epochs as $80$, GMem enhances the sampling efficiency by 10×10$\times$10 ×, where GMem requires only $25$ sampling steps to achieve an FID $=12.3$ while the vanilla SiT needs $250$ sampling steps, yielding 10×10$\times$10 × sampling time savings.
We outline the main contributions of this paper below:
@@
    We propose a novel conjecture that interprets the functionality of diffusion models into two distinct components: (i) memorization of semantic information and (ii) generalization to the true data distribution.
@@
    We introduce GMem, a simple yet highly effective approach that decouples the memorization capability of diffusion models into an external memory bank. This separation alleviates the memorization burden on the neural networks and enhances generalization flexibility.
@@
    GMem significantly enhances efficiency and quality of the SoTA diffusion models, for both training and inference (see Figure 1 and Figure 2).
Figure 2: Selected samples on ImageNet $512\times 512$ and $256\times 256$. This figure presents images generated by GMem under two experimental settings: (1) For ImageNet $256\times 256$, GMem was trained for $160$ epochs and sampled via Euler method (NFE $=100$), achieving an FID $=1.53$ without classifier-free guidance. (2) For ImageNet $512\times 512$, training extended to $400$ epochs with identical sampling settings, yielding FID $=1.89$.
@@ 2 Related work
Generative models–—including Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Sauer et al., 2022; Xiao et al., 2021), Variational Autoencoders (VAEs) (Kingma, 2013; He et al., 2022), flow-based methods, and diffusion-based methods (Ho et al., 2020; Dhariwal & Nichol, 2021; Mittal et al., 2023)—–aim to learn the data distribution $p(\mathbf{x})$ and generate data through sampling, achieving remarkable performance in producing realistic samples (Li et al., 2023d). Recently, diffusion-based methods employ stochastic interpolation to model a forward process and then reverse the Gaussian distribution back to the original image space, generating realistic samples. These methods achieve SoTA results in deep generative modeling and are the focus of this study (Mittal et al., 2023; Song & Ermon, 2020; Durkan & Song, 2021).
Diffusion models face computational challenges due to high training cost and instability (Yu et al., 2024; Song & Ermon, 2020) and high sampling costs from multi-step generation (Lu & Song, 2024), driving extensive research to accelerate both processes. For example, REPA (Yu et al., 2024) leverages external visual representations to speed up training. LightningDiT (Yao & Wang, 2025) accelerates training by aligning the latent space of vision tokenizer (i.e. VA-VAE) with pretrained vision encoder. GMem instead introduces a novel decomposition of diffusion models into generalization and memorization components, enabling significantly faster training.
@@
To overcome the instability and computational inefficiency of diffusion models, recent studies (Yu et al., 2024; Fuest et al., 2024; Mittal et al., 2023) start to leverage representation learning to enhance diffusion models. On the one hand, diffusion models are capable of learning high-quality representations (Yu et al., 2024). For instance, Tang et al. (2023) demonstrate that feature maps extracted from diffusion networks can establish correspondences between real images, indicating a strong alignment between the learned representations with actual image. Furthermore, Yang & Wang (2023) conduct a detailed analysis of the trade-off between the quality of learned representations and the penalization of the optimal parameter spectrum.
@@
On the other hand, well-trained representation models can improve performance and expedite the training of diffusion models. Mittal et al. (2023) accomplish this by adjusting the weighting function in the denoising score matching objective to enhance representation learning. REPA (Yu et al., 2024) introduces an alignment loss for intermediate layer representations, significantly accelerating the training process by over $17.5$ times.
External representation-augmented diffusion models via retrieval and generation.
Retrieval-Augmented Generation (RAG) enhances generation quality by integrating external knowledge. IC-GAN (Casanova et al., 2021) augments image generation by conditioning on neighborhood instances retrieved from the training dataset. However, using only training images limits generalization. To address this issue, KNN-Diffusion (Sheynin et al., 2022) and RDM (Blattmann et al., 2022) employ large external memory sets, guiding generation via kNN retrieval during training and inference. Similarly, Chen et al. (2022) and Li et al. (2022) leverage a set of text-image pairs with cross-modality retrieval, improving generation performance on rare images.
Despite their advantages, RAG methods encounter two key challenges: (1) substantial storage demands for large memory sets, and (2) increased computational costs during retrieval. We further contend that over-reliance on training sets restricts generalization capabilities (Blattmann et al., 2022). By employing a masking strategy (see our Section 5.4), we mitigate this dependency without incurring additional computational or storage overhead, thereby improving both generalization and training/inference efficiency.
Representations for generation augmentation can also be obtained from representation generators. For instance, RCG (Li et al., 2023b) employs a representation generator to produce 1D “memory snippets” to guide diffusion models. Although RCG reduces the need to store large-scale memory sets, it encounters two primary challenges: (1) the requirement for additional training and sampling processes for the representation generator, increasing computational demands; (2) the necessity to retrain the representation generator to incorporate knowledge of new classes or artistic style transfers, thereby limiting its generalization capability. To overcome these limitations, Section 5.3 presents an efficient, training-free method for incorporating additional knowledge into the memory bank.
@@ 3 Preliminary
In this section, we provide a concise introduction to the training and sampling processes of flow-based and diffusion-based models. See Appendix F for more details.
Both diffusion-based (Ho et al., 2020; Dhariwal & Nichol, 2021) and flow-based models (Ma et al., 2024) derive their training procedures from a deterministic $T$-step noising process applied to the original data (Ma et al., 2024), formalized as:
@@
According to (1), each marginal probability density $p_{t}(\mathbf{x}_{t})$ corresponds to the distribution of a Probability Flow Ordinary Differential Equation (Song et al., 2020b) (PF ODE) with velocity field $\mathbf{v}(\mathbf{x},t)$, defined as:
@@
Alternatively, the aforementioned noise-adding process can be formalized as a Stochastic Differential Equation (Song et al., 2020b, a) (SDE):
@@
where $\mathbf{W}_{t}$ is a Wiener process (Hitsuda, 1968), $\boldsymbol{m}(\mathbf{x}_{t},t)$ is the drift coefficient defined as $\boldsymbol{m}(\mathbf{x}_{t},t)=-\frac{1}{2}\beta(t)\mathbf{x}_{t}$, and $g(t)$ is the diffusion coefficient, set as $g(t)=\sqrt{\beta(t)}$ with $\beta(t)$ being a time-dependent positive function controlling the noise schedule.
@@
By solving the reverse-time SDE in (4), starting from the initial state $\mathbf{x}_{T}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, we can obtain $p_{0}(\mathbf{x}_{0})$, thereby estimating the true data distribution $p(\mathbf{x})$.
@@ 4 Methodology
[…]
GMem is motivated by the observation that diffusion models inherently encode two distinct capabilities: generalization and memorization of semantic information within data (Yu et al., 2024; Yang & Wang, 2023; Xiang et al., 2023; Chen et al., 2024c). We introduce a novel paradigm that decouples these functions: neural networks handle generalization, while an external memory bank stores semantic information.
Current generative models (Ho et al., 2020; Song & Ermon, 2020) typically rely on one neural network to simultaneously achieve both generalization and memorization of data distributions. However, the capacity-constrained modern diffusion models, such as UNet and Transformer-based networks (Ho et al., 2020; Peebles & Xie, 2023), face two critical limitations (Kadkhodaie et al., 2023): (1) insufficient model parameters to memorize complex data distributions, and (2) computationally expensive parameter optimization for memorization.
To overcome these limitations, we propose offloading the memorization function to an external memory bank, thereby enhancing training efficiency and distribution capture capabilities. The proposed framework is illustrated in Figure 3.